# On Sybil-Proof Mechanisms

### v4

@@ Abstract
We show that in the single-parameter mechanism design environment, the only non-wasteful, symmetric, incentive compatible and Sybil-proof direct mechanism is a second price auction with symmetric tie-breaking. Thus, if there is private information, lotteries or other mechanisms that do not always allocate to a highest-value bidder are not Sybil-proof or not incentive compatible. Moreover, we show that our main (im)possibility result extends beyond linear valuations, but not to multi-unit object allocation with capacity constrained bidders.
@@ 1 Introduction
In the case of online sales of event tickets, tickets are often sold at a relatively low price where there is still excess demand. The “under-pricing” of tickets is a way for the event organizers to give dedicated fans with smaller budgets a chance to participate, yet this also makes the primary sale a target for middlemen that resell tickets with a premium on a secondary market Budish and Bhave, (2023). If quantities of tickets per buyer are capped, middlemen effectively achieve this by using Sybils in the primary sale. [^1] Related issues also appear in crypto-currency “airdrops” where protocols want to reward early adopters with tokens, see Messias et al., (2023), and the use of non-proportional rules, has lead to wide-spread exploits through Sybils who afterwards sell their airdrop on a secondary market.
@@
As another example, in blockchain systems, randomness in allocating the right to propose blocks is desirable to avoid power concentration, which could lead to censorship and undermine decentralization. Bitcoin’s proof-of-work mechanism in particular, which uses pseudorandomness in proposer selection, has been argued to be an effective method to achieve decentralization of the system by having many different actors “mining” blocks. [^2] The proof-of-work mechanism of Bitcoin is a particular instance of a proposer selection mechanism that assign the right to propose the next block with probability proportional to effort (of mining). Similarly, in proof-of-stake systems, proposal rights are usually assigned through a lottery with chances proportionally to stake. Proportional selection mechanisms have in this context been characterized by Sybil-proofness, Symmetry, Non-wastefulness, and Collusion-proofness Chen et al., (2019); Leshno and Strack, (2020). Under this mechanism, one’s chance of winning depends, however, on others’ effort. Thus, the mechanism fails to be incentive compatible. This assessment can, however, change if there is significant heterogeneity in value for proposal rights. [^3] Heterogeneous (and private) value is a concern, as different participants can generate different values from block proposal rights. For example, some block proposers have exclusive access to some of the submitted transactions to include in their block, as documented in Öz et al., (2024) for the case of Ethereum. In practice, this had led to the creation of out-of-protocol secondary market for the content of blocks, and these markets exhibit a high degree of market concentration in the hand of few block builders Öz et al., (2024); in other words, empirically the assertion of a “monopoly without monopolist” Huberman et al., (2021) does not really hold.
@@
Given these considerations, it is natural to ask whether there are Sybil-proof mechanisms for assigning private goods other than auctions. Our main result, Theorem 1, gives a strong negative answer to this question: in the classical Myersonian mechanism design setting with quasi-linear preferences and one-dimensional types, we show that the only monotonic, symmetric, and non-wasteful allocation rule for which the induced mechanism with “Myerson payments” is Sybil-proof, is the rule that allocates everything to the highest value bidder, breaking ties uniformly (if the good is indivisible), resp. sharing the unit equally among highest value bidders (if the good is divisible). Thus, while there is a large design space of monotonic allocation rules - as two extreme cases we could assign the good with equal probability, independently of value, among participants or we could always assign the good to the highest value bidder - adding Sybil-proofness collapses the set of implementable allocation rules to a single one. The result extends beyond linear valuation (Theorem 3), but interestingly not to the multi-unit case with unit-demand bidders (Proposition 1).
@@
While our main theorem is formulated for direct mechanisms that elicit values from the participants, indirect mechanism cannot circumvent the impossibility, as we have a version of the revelation principle in the presence of Sybils: for every symmetric indirect mechanism with Sybils that implements an objective function in dominant strategies, the direct mechanism induced by the objective function is Sybil-proof and incentive compatible.
@@ 1.1 Related Work and Implications
Sybil-proof mechanisms have been studied in various contexts, including voting systems Wagman and Conitzer, (2008), combinatorial auctions Yokoo et al., (2004); Gafni et al., (2020); Gafni and Tennenholtz, (2023), recommender systems Brill et al., (2016) and blockchain systems Chen et al., (2019); Leshno and Strack, (2020), and different Sybil-proof mechanisms [^4] Previously literature also used the term false-name proof mechanisms for a stronger notion than the one we use in this paper. have been discussed in these contexts. Most related to our work, Yokoo et al., (2004) prove that there is no Pareto-efficient, false-name-proof combinatorial auction. We do not require Pareto efficiency, as we are primarily interested in whether non-trivial distributional goals can be achieved. However, it is as a consequence of our theorem that the combination of Non-wastefulness, Sybil-proofness and Incentive Compatibility implies Pareto-efficiency.
@@
Our results also relate to the literature on mechanism design with one-dimensional types. In the case without Sybils, for one-dimensional types Dominant-Strategy- and Bayesian- incentive compatibility are equivalent in the sense that for each Bayes-Nash IC mechanism one can construct a Dominant-Strategy IC mechanism that gives each participant the same interim expected pay-off Manelli and Vincent, (2010); Gershkov et al., (2013). We show that the statement is no-longer true with additional Sybil-proofness constraints and dominant-strategy implementation is strictly stronger than Bayes-Nash implementation in the presence of Sybils.
@@
Similar but different impossibility results also occur if one considers collusion resistance instead of Sybil resistance as a desideratum. [^5] Collusion differs from sybilling in so far as different agents colluding, in general, have different preferences over outcomes of a mechanism. Moreover, collusion is usually studied in a fixed population setting, whereas sybilling means creating (an arbitrary number of) new participants with alligned preferences. For transaction fee mechanisms for blockchains for example, Gafni and Yaish, (2024) show that there is no non-trivial mechanism that satisfies constant probability allocation, incentive compatibility, miner incentive compatibility and collusion resistance.
2 Model and Result
@@
We consider mechanisms with a variable population of bidders: A (direct) mechanism specifies for each finite $\mathcal{N}\subset\mathbb{N}$ set of bidders (some of them real agents, some of them, possibly, Sybils), an allocation rule $x^{\mathcal{N}}:\mathbb{R}_{+}^{\mathcal{N}}\rightarrow\Delta(\mathcal{N})$ where $\Delta(\mathcal{N}):=\{x\in\mathbb{R}_{+}^{\mathcal{N}}:\sum_{i\in\mathcal{N}}x_{i}\leq 1\}$ and a payment rule $p^{\mathcal{N}}:\mathbb{R}^{\mathcal{N}}_{+}\rightarrow\mathbb{R}^{\mathcal{N}}.$ In the following, we will often omit the superscript $\mathcal{N}$ when there is no ambiguity. We can interpret the allocation shares $x^{\mathcal{N}}_{i}(v)$ for the reported values [^6] We work with the non-negative reals as type space. Alternatively, we could also work with an interval $0,\bar{v}]$ as type space and obtain the same characterization result with completely analogous proofs. $v\in\mathbb{R}_{+}^{\mathcal{N}}$ either as probabilities of obtaining an indivisible good or as an allocation of a perfectly divisible good of which one unit is distributed in total. A bidder $i$ with value $v_{i}\geq 0$ has a linear utility
$U_{i}(x_{i},p_{i})=v_{i}x_{i}-p_{i}$
@@
Next, we introduce several axioms that the mechanisms should satisfy. First we want participation to be voluntary, in the sense that bidders who do not derive value from the item do not need to pay:
@@
As known from classical results Myerson, (1981), this is equivalent to using “Myerson payments”,
@@
Fifth, we want the mechanism to be Sybil-proof. Previous literature has used the term “false-name-proofness”, Yokoo et al., (2004), but usually for the combination of incentive compatibility and Sybil-proofness, which requires immunity to deviations where the bidder reports a different value and creates Sybils. For our result a weaker notion of Sybil-proofness is needed, which only requires immunity to Sybil attacks where a bidder reports truthfully from his original account and creates one Sybil with an arbitrary bid:
@@
2.1 The Revelation Principle for Sybils
In principle, it might not be bad in itself that participants in a mechanism use Sybils, as long as the mechanism achieves good outcomes. However, it turns out that the possibility to sybil fundamentally constrains what outcomes can be achieved by any mechanism. This is a variant of the “revelation principle”. If an arbitrary symmetric indirect mechanism implements an objective in such a way that bidders might use Sybils as part of their strategies, then there is necessarily a corresponding direct mechanism achieving the same objective in a Sybil-proof way. For the technical definition of a symmetric indirect mechanism, its extensions with Sybils, the notion of ”implementation” and the Revelation Principle see the Appendix. The revelation principle allows us to focus on direct mechanisms that are IC and Sybil-Proof subsequently.
2.2 Characterization Result
@@
By Myerson’s lemma, the payments that implement the rule in dominant strategies are defined by Equation (1). Subsequently, we will use the following two facts about the payments: first payment normalization implies individual rationality under Myerson payments
@@
and the payment rule given by the “Myerson payments”. Then the mechanism satisfies Symmetry, Incentive Compatibility, Monotonicity, and Sybil-proofness. The only nontrivial property to check is Sybil-proofness which we establish in the appendix.
@@ 3 Bayesian Sybil-Proofness
Technically, for the Bayesian version of our model we need to enhance it by a prior which is a probability distribution $\mu$ over $\bigcup_{\mathcal{N}\subset\mathbb{N},|\mathcal{N}|<\infty}\mathbb{R}_{+}^{\mathcal{N}}$. Thus, both, valuations and the number of real bidders, can be unknown. Denote by $\mu(\cdot|v_{i})$ the posterior of bidder $i$ if he is part of the mechanism and has type $v_{i}$. The posterior is a probability distribution over $\bigcup_{i\in\mathcal{N}\subset\mathbb{N}}\mathbb{R}_{+}^{\mathcal{N}}.$ Subsequently, we focus on the independent private value setting, where the conditional distribution of values, conditional on the set of agents in the mechanism being $\mathcal{N}$, is $F^{\mathcal{N}}$ for a distribution $F$ on $\mathbb{R}_{+}$ that is twice differentiable and strictly increasing on its support (assumed to be $(0,\bar{v})$ for some $\bar{v}>0$ or $\bar{v}=\infty$) with $F(0)=0$.
@@
Moreover, we want the mechanism to be Bayesian Sybil-proof. We consider a rather strong Bayesian notion of Sybil-proofness, which is similar to the one in Gafni et al., (2020); Mazorra and Della Penna, (2023). The notion requires that creating Sybils does not benefit a bidder in expectation if no-one else uses a Sybil in equilibrium and reports their true valuation.
@@
We consider a lottery mechanism where the item is allocated with equal chances among agents that pay an entry price and the ”ticket price” is increasing in the number of participants.
@@
which we call $a_{n}$ the ”ticket price”. The lottery mechanism with increasing ticket price for $n$ bids is defined as follows:
@@
Suppose agents have valuations drawn from a distribution $F$ strictly increasing with support $[0,1]$. When all agents report truthfully, any agent $i$’s ex-interim expected payoff (conditional on knowing the number of submitted bids) under the lottery with increasing ticket price mechanism is decreasing in the ticket price $a\in[y^{\star},1]$. Moreover, any agent’s ex-interim expected payoff is at least as high as in a standard second-price auction.
@@
We establish the result for the case that the number of bidders is known by the bidders. A fortiori this also establishes the result if the number of bidders follows a non-degenerate prior. If $v<a_{n}$, the agent cannot profit by splitting into multiple identities: the mechanism behaves like a second-price auction, where multi‐bid strategies do not help since the second-price auction is Sybil-proof. Hence assume $v\geq a_{n}$.
Suppose the agent creates $k=k_{1}+k_{2}$ Sybils, with $k_{1}$ Sybils bidding a value greater or equal $a_{n-1+k}$ and $k_{2}$ Sybils bidding a value below $a_{n-1+k}$. Because $\{a_{n}\}$ is increasing and $a_{n}\geq y^{\star}$, Lemma 4 rules out profitable manipulations when $k_{1}=1$ and $k_{2}\geq 2$.
@@
A monotonicity argument, see Lemma 6 in the Appendix, shows $U(v,k_{1},k_{2},a)$ is monotone non‐increasing in $a\in[F^{-1}(\frac{n-2}{n-1}),1]$; thus
$U(v,k_{1},k_{2},a_{n-1+k})\;\leq\;U(v,k_{1},k_{2},a_{n-1+k_{1}})\;\leq\;v-a_{n-1+k_{1}}.$
On the other hand,
$v-a_{n-1+k_{1}}$ $\leq v-a_{n}$
$\leq(v-E[Z_{n-1}\mid Z_{n-1}\leq v])\Pr[Z_{n-1}\leq v]$
$\leq U(v,a_{n})$
The first inequality is deduced from the fact that the sequence $a_{n}$ is increasing. The third inequality is deduced by Lemma 4 since $(v-E[Z_{n-1}\mid Z_{n-1}\leq v])\Pr[Z_{n-1}\leq v]$ is the ex-interim expected payoff in the second-price auction. The second inequality holds since $v-a_{n}\leq(v-E[Z_{n-1}\mid Z_{n-1}\leq v])\Pr[Z_{n-1}\leq v]$ if and only if $v-(v-E[Z_{n-1}\mid Z_{n-1}\leq v])\Pr[Z_{n-1}\leq v]\leq a_{n}$. This last inequality is deduced from
$v-(v-E[Z_{n-1}\mid Z_{n-1}\leq v])\Pr[Z_{n-1}\leq v]$ $=v\Pr[Z_{n-1}>v]+E[Z_{n-1}\mid Z_{n-1}\leq v]\Pr[Z_{n-1}\leq v]$
$\leq E[Z_{n-1}]$
$\leq a_{n}\text{ (By construction)}.$
Hence Sybil strategies yield lower expected payoffs, proving the mechanism is Bayesian Sybil‐proof. ∎

### v5

@@ Abstract
We show that in the single-parameter mechanism design environment, the only non-wasteful, symmetric, incentive compatible and Sybil-proof direct mechanism is a second price auction with symmetric tie-breaking. Thus, if there is private information, lotteries or other mechanisms that do not always allocate to a highest-value bidder are not Sybil-proof or not incentive compatible. Moreover, we show that our main (im)possibility result extends beyond linear valuations, but not to multi-unit object allocation with unit-demand bidders.
@@ 1 Introduction
In the case of online sales of event tickets, tickets are often sold at a relatively low price where there is still excess demand. The “under-pricing” of tickets is a way for the event organizers to give dedicated fans with smaller budgets a chance to participate, yet this also makes the primary sale a target for middlemen that resell tickets with a premium on a secondary market Budish and Bhave (2023). If quantities of tickets per buyer are capped, middlemen effectively achieve this by using Sybils in the primary sale. [^1] Related issues also appear in crypto-currency “airdrops” where protocols want to reward early adopters with tokens, see Messias et al. (2023), and the use of non-proportional rules, has lead to wide-spread exploits through Sybils who afterwards sell their airdrop on a secondary market.
@@
As another example, in blockchain systems, randomness in allocating the right to propose blocks is desirable to avoid power concentration, which could lead to censorship and undermine decentralization. Bitcoin’s proof-of-work mechanism in particular, which uses pseudorandomness in proposer selection, has been argued to be an effective and Sybil-proof method to achieve decentralization of the system by having many different actors “mining” blocks. This assessment can, however, change if there is significant heterogeneity in value for proposal rights. [^2] Heterogeneous (and private) value is a concern, as different participants can generate different values from block proposal rights. For example, some block proposers have exclusive access to some of the submitted transactions to include in their block, as documented in Öz et al. (2024) for the case of Ethereum. In practice, this had led to the creation of out-of-protocol secondary market for the content of blocks, and these markets exhibit a high degree of market concentration in the hand of few block builders Öz et al. (2024); in other words, empirically the assertion of a “monopoly without monopolist” Huberman et al. (2021) does not really hold.
@@
Given these considerations, it is natural to ask whether there are Sybil-proof mechanisms for assigning private goods other than auctions. Our main result, Theorem 1, gives a strong negative answer to this question: in the classical Myersonian mechanism design setting with quasi-linear preferences and one-dimensional types, we show that the only monotonic, symmetric, and non-wasteful allocation rule for which the induced mechanism with “Myerson payments” is Sybil-proof, is the rule that allocates everything to the highest value bidder, breaking ties uniformly (if the good is indivisible), resp. sharing the unit equally among highest value bidders (if the good is divisible). Thus, while there is a large design space of monotonic allocation rules - as two extreme cases we could assign the good with equal probability, independently of value, among participants or we could always assign the good to the highest value bidder - adding Sybil-proofness collapses the set of implementable allocation rules to a single one. The result extends beyond linear valuation (Proposition 4), but interestingly not to the multi-unit case with unit-demand bidders (Proposition 3).
@@
While our main theorem is formulated for direct mechanisms that elicit values from the participants, indirect mechanism cannot circumvent the impossibility, as we have a version of the revelation principle in the presence of Sybils (Section 5): for every symmetric indirect mechanism with Sybils that implements an objective function in dominant strategies, the direct mechanism induced by the objective function is Sybil-proof and incentive compatible.
@@ 1.1 Related Work
Sybil-proof mechanisms have been studied in various contexts, including voting systems Wagman and Conitzer (2008), combinatorial auctions Yokoo et al. (2004); Gafni et al. (2020); Gafni and Tennenholtz (2023), recommender systems Brill et al. (2016) and blockchain systems Chen et al. (2019); Leshno and Strack (2020), and different Sybil-proof mechanisms [^3] Previously literature also used the term false-name proof mechanisms for a stronger notion than the one we use. have been discussed in these contexts. Most related to our work, Yokoo et al. (2004) prove that there is no Pareto-efficient, false-name-proof combinatorial auction. We do not require Pareto efficiency, as we are primarily interested in whether non-trivial distributional goals can be achieved. However, it is as a consequence of our theorem that the combination of Non-wastefulness, Sybil-proofness and Incentive Compatibility implies Pareto-efficiency.
@@
Our results also relate to the literature on mechanism design with one-dimensional types. In the case without Sybils, for one-dimensional types Dominant-Strategy- and Bayesian- incentive compatibility are equivalent in the sense that for each Bayes-Nash IC mechanism one can construct a Dominant-Strategy IC mechanism that gives each participant the same interim expected pay-off Manelli and Vincent (2010); Gershkov et al. (2013). We show that the statement is no-longer true with additional Sybil-proofness constraints and dominant-strategy implementation is strictly stronger than Bayes-Nash implementation in the presence of Sybils.
@@
Similar but different impossibility results also occur if one considers collusion resistance instead of Sybil resistance as a desideratum. [^4] Collusion differs from sybilling in so far as different agents colluding, in general, have different preferences over outcomes of a mechanism. Moreover, collusion is usually studied in a fixed population setting, whereas sybilling means creating (an arbitrary number of) new participants with alligned preferences. For transaction fee mechanisms for blockchains for example, Gafni and Yaish (2024) show that there is no non-trivial mechanism that satisfies constant probability allocation, incentive compatibility, miner incentive compatibility and collusion resistance.
2 Model
@@
We consider mechanisms with a variable population of bidders: A (direct) mechanism  [^5] Focusing on direct mechanism, where participants report their value to the mechanism which determines an allocation from the reports is without loss of generality by the revelation principle, which we formally establish in Section 5. specifies for each finite $\mathcal{N}\subset\mathbb{N}$ set of bidders (some of them real agents, some of them, possibly, Sybils), an allocation rule $x^{\mathcal{N}}:\mathbb{R}_{+}^{\mathcal{N}}\rightarrow\Delta(\mathcal{N})$ where $\Delta(\mathcal{N}):=\{x\in\mathbb{R}_{+}^{\mathcal{N}}:\sum_{i\in\mathcal{N}}x_{i}\leq 1\}$ and a payment rule $p^{\mathcal{N}}:\mathbb{R}^{\mathcal{N}}_{+}\rightarrow\mathbb{R}^{\mathcal{N}}.$ In the following, we will often omit the superscript $\mathcal{N}$ when there is no ambiguity. We can interpret the allocation shares $x^{\mathcal{N}}_{i}(v)$ for the reported values [^6] We work with the non-negative reals as type space. Alternatively, we could also work with an interval $0,\bar{v}]$ as type space and obtain the same characterization result with completely analogous proofs. $v\in\mathbb{R}_{+}^{\mathcal{N}}$ either as probabilities of obtaining an indivisible good or as an allocation of a perfectly divisible good of which one unit is distributed in total. A bidder $i$ with value $v_{i}\geq 0$ has a linear utility
$U_{i}=v_{i}x_{i}-p_{i}$
@@
Next, we introduce several axioms that the mechanisms should satisfy. First, we want participation to be voluntary, in the sense that bidders who do not derive value from the item do not need to pay:
@@
As known from classical results Myerson (1981), if zero value bidders do not pay, this is equivalent to using “Myerson payments”,
@@
Fifth, we want the mechanism to be Sybil-proof. Previous literature has used the term “false-name-proofness”, Yokoo et al. (2004), but usually for the combination of incentive compatibility and Sybil-proofness, which requires immunity to deviations where the bidder reports a different value and creates Sybils. For our result a weaker notion of Sybil-proofness is needed, which only requires immunity to Sybil attacks where a bidder reports truthfully from his original account and creates one Sybil with an arbitrary bid:
@@
Remark 1.
There are other closely related notions of (ex-post) Sybil-proofness that we could use instead and still obtain our main result. Note that, as we require immunity to manipulations through the creation of one Sybil, the following main result will hold a forteriori also for stronger notions of Sybil-proofness that require immunity to manipulation through the creation of an arbitrary number of Sybils. In particular, our main result holds with Sybil-Proofness (and incentive compatibility) replaced by the False-Name-Proofness axiom studied in the previous literature. However, we use our version of the Sybil-proofness axiom to establish an impossibility result with as weak of an ex-post Sybil-proofness axiom as possible yielding a stronger result.
Another natural variant of Sybil-proofness would be to require that for each bid vector representing multiple bids of one agent, there is a single bid that weakly dominates the Sybil-bid. With this notion, our main result would again hold true, as Incentive Compatibility and this version of Sybil-Proofness imply again False-Name-Proofness which implies our version of Sybil-Proofness (and Incentive Compatibility).
3 Characterization Result
@@
Note that payment normalization and Sybil-proofness imply that $p_{j}(0,v_{-j})=0$ for a bidder $j$ who bids 0. Thus, by Myerson’s lemma, the payments that implement the rule in dominant strategies are defined by Equation (1). Subsequently, we will use the following two facts about the payments: first we have individual rationality under Myerson payments
@@
and the payment rule given by the “Myerson payments”.Then the mechanism satisfies Payment Normalization, Symmetry, Incentive Compatibility, Monotonicity, and Sybil-proofness. The only nontrivial property to check is Sybil-proofness.
Suppose that there are $N$ bidders with values $u$. If they all bid truthfully and we define $C:=\sum_{j>1}u_{j}$, Bidder $1$ has utility payoff,
$U_{1}=2^{-N}\int_{0}^{u}\frac{z}{z+C}dz$
Suppose that Bidder $1$ deviates by creating a Sybil bidding $x$. Then his utility payoff will be
$2^{(N+1)}U^{\prime}_{1}$ $=$ $\int_{0}^{u}\frac{z}{z+x+C}dz+\frac{x(u_{1}-x)}{u_{1}+x+C}+\int_{0}^{x}\frac{z}{z+u_{1}+C}dz$
We show that $U_{1}^{\prime}\leq U_{1}$ for any $x\geq 0$, i.e. that bidder $1$ has no incentive to deviate by creating Sybils. We have
$2^{N+1}(U_{1}-U_{1}^{\prime})$ $=2\int_{0}^{u_{1}}\frac{z}{z+C}dz-\int_{0}^{u_{1}}\frac{z}{z+x+C}dz-\frac{x(u_{1}-x)}{u_{1}+x+C}-\int_{0}^{x}\frac{z}{z+u_{1}+C}dz$
$\geq\int_{0}^{u_{1}}\frac{z}{z+C}dz-\frac{x(u_{1}-x)}{u_{1}+x+C}-\int_{0}^{x}\frac{z}{z+u_{1}+C}dz$
Note that the expression on the right hand side of the inequality is positive on the boundary i.e. at $x=0$ and at $x=\infty$ (as it diverges to $\infty$). The expression on the right hand side of the inequality has one critical point at $x=u_{1}$, and the expression is positive at that critical point. Therefore the difference in utility is bounded from below with a strictly positive expression.
@@ 4 Bayesian Sybil-Proofness
Technically, for the Bayesian version of our model we need to enhance it by a prior which is a probability distribution $\mu$ over $\bigcup_{\mathcal{N}\subset\mathbb{N},|\mathcal{N}|<\infty}\mathbb{R}_{+}^{\mathcal{N}}$. Thus, both, valuations and the number of real bidders, can be unknown. Denote by $\mu(\cdot|v_{i})$ the posterior of bidder $i$ if he is part of the mechanism and has type $v_{i}$. The posterior is a probability distribution over $\bigcup_{i\in\mathcal{N}\subset\mathbb{N}}\mathbb{R}_{+}^{\mathcal{N}}.$ Subsequently, we focus on the independent private value setting, where the conditional distribution of values, conditional on the set of agents in the mechanism being $\mathcal{N}$, is $F^{\mathcal{N}}$ for a distribution $F$ on $\mathbb{R}_{+}$ that is twice differentiable and strictly increasing on its support (assumed to be $(0,\bar{v})$ for some $\bar{v}>0$) with $F(0)=0$.
@@
Moreover, we want the mechanism to be Bayesian Sybil-proof. We consider a rather strong Bayesian notion of Sybil-proofness, which is similar to the one in Gafni et al. (2020); Mazorra and Della Penna (2023). The notion requires that creating Sybils does not benefit a bidder in expectation if no-one else uses a Sybil in equilibrium and reports their true valuation.
@@
We consider a lottery mechanism where the item is allocated with equal chances among agents that pay an entry price and the “ticket price” is increasing in the number of participants.
@@
which we call $a_{n}$ the “ticket price”. The lottery mechanism with increasing ticket price for $n$ bids is defined as follows:
@@
Suppose agents have valuations drawn from a distribution $F$ strictly increasing with support $[0,1]$. When all agents report truthfully, any agent $i$’s ex-interim expected payoff conditional on knowing the number of submitted bids under the lottery with increasing ticket price mechanism is decreasing in the ticket price $a\in[y^{\star},1]$. Moreover, any agent’s ex-interim expected payoff is at least as high as in a standard second-price auction.
@@
We establish the result for the case that the number of bidders is known by the bidders. A fortiori this also establishes the result if the number of bidders follows a non-degenerate prior. An agent with value $v<a_{n}$ cannot profit by splitting into multiple identities: the mechanism behaves like a second-price auction, where multi‐bid strategies do not help since the second-price auction is Sybil-proof. Hence consider an agent with value $v\geq a_{n}$.
Suppose the agent creates $k=k_{1}+k_{2}$ Sybils, [^10] Subsequently, we use the term “creates $k$ Sybils” for the case where the agent bids from $k$ accounts (his “original” account and k1k-1 “fake” accounts). with $k_{1}$ Sybils bidding a value greater or equal $a_{n-1+k}$ and $k_{2}$ Sybils bidding a value below $a_{n-1+k}$. Because $\{a_{n}\}$ is increasing and $a_{n}\geq y^{\star}$, Lemma 4 rules out profitable manipulations when $k_{1}=1$ and $k_{2}\geq 2$.
@@
Since $0\leq\frac{k_{1}}{m+k_{1}}\leq 1$ for every $m\in\{0,\dots,n-1\}$, we have
$0\leq\sum_{m=0}^{n-1}\binom{n-1}{m}\bigl(1-F(a_{n-1+k})\bigr)^{m}F(a_{n-1+k})^{\,n-1-m}\frac{k_{1}}{m+k_{1}}\leq 1.$
If $a_{n-1+k}>v$, then $U(v,k_{1},k_{2},a_{n-1+k})\leq 0$. Otherwise, if $a_{n-1+k}\leq v$, then
$U(v,k_{1},k_{2},a_{n-1+k})$ $\leq(v-a_{n-1+k})\sum_{m=0}^{n-1}\binom{n-1}{m}\bigl(1-F(a_{n-1+k})\bigr)^{m}F(a_{n-1+k})^{\,n-1-m}$
$=v-a_{n-1+k},$
since the binomial weights sum to $1$. Hence, in either case,
$U(v,k_{1},k_{2},a_{n-1+k})\leq\max\{v-a_{n-1+k},0\}\leq v-a_{n},$
where the last inequality uses that $\{a_{t}\}$ is increasing (so $a_{n-1+k}\geq a_{n}$) and that we are in the case $v\geq a_{n}$. On the other hand,
$v-a_{n}$ $\leq(v-\mathbb{E}[Z_{n-1}\mid Z_{n-1}\leq v])\,\Pr[Z_{n-1}\leq v]$
$\leq U(v,a_{n}),$
where $U(v,a_{n})$ denotes the interim payoff under the lottery mechanism with ticket price $a_{n}$ for the bidder with valuation $v$. The second inequality is deduced by Lemma 4 since $(v-\mathbb{E}[Z_{n-1}\mid Z_{n-1}\leq v])\,\Pr[Z_{n-1}\leq v]$ is the ex-interim expected payoff in the second-price auction. The first inequality holds since
$v-(v-\mathbb{E}[Z_{n-1}\mid Z_{n-1}\leq v])\,\Pr[Z_{n-1}\leq v]$ $=v\,\Pr[Z_{n-1}>v]+\mathbb{E}[Z_{n-1}\mid Z_{n-1}\leq v]\,\Pr[Z_{n-1}\leq v]$
$\leq\mathbb{E}[Z_{n-1}]\leq a_{n},$
where the last inequality is by construction of $a_{n}$.
Combining the two displays yields $U(v,k_{1},k_{2},a_{n-1+k})\leq U(v,a_{n})$. Hence Sybil strategies yield lower expected payoffs, proving the mechanism is Bayesian Sybil‐proof. ∎
5 The Revelation Principle for Sybils
In principle, it might not be bad in itself that participants in a mechanism use Sybils, as long as the mechanism achieves good outcomes. However, it turns out that the possibility to sybil fundamentally constrains what outcomes can be achieved by any mechanism. This is a variant of the “revelation principle”. If an arbitrary symmetric indirect mechanism implements an objective in such a way that bidders might use Sybils as part of their strategies, then there is necessarily a corresponding direct mechanism achieving the same objective in a Sybil-proof way. We formally establish this revelation principle in this section, thereby justifying our focus on direct mechanisms in the previous sections.
Generally, subtleties arise around the players reasoning about the true number of bidders. Subsequently, we assume that bidders cannot condition their message on the number of messages the Sybil mechanism receives (because they do not know it when sending the message to the mechanism). For the dominant strategy implementation case, we can then operate in a “belief-free” model: sending message $m_{i}(v_{i})$ should be weakly-dominant for a bidder $i$ with value $v_{i}$, for all possible sets of participants in the Sybil mechanism. For the Bayesian setting bidders will also hold beliefs about the number (or more generally set of) bidders that participate in the Sybil mechanism, as we further discuss below.
A social welfare function for a variable population model is a function $f:\mathcal{D}\to\Delta(\mathbb{N})\times\mathbb{R}_{+}^{\mathbb{N}}$ with $f_{i}(v_{1},\ldots)=(0,0)$ for $v_{i}=0$ where $\mathcal{D}\subseteq c_{00}(\mathbb{R}_{+})$ is a subset of the set of $\mathbb{R}_{+}$-valued sequences with finitely many non-zero elements. The first element of $f_{i}(v)$ corresponds to the probability with which bidder $i$ gets the item, and the second element corresponds to the payment bidder $i$ makes. A social welfare function yields a direct mechanism where $\mathcal{M}=\mathbb{R}_{+}$ and $(x^{\mathcal{N}}(v),p^{\mathcal{N}}(v))=f(v|_{\mathcal{N}},0|_{\mathbb{N}\setminus\mathcal{N}})|_{\mathcal{N}}$.
[…]
  1. 1.
[…]
  2. 2.
We have the following proposition which is similar to an observation by Yokoo et al. (2004):
[…]
Proof.
[…]
  1. 1.
  2. 2.
[…]
Proof.